40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Dec 17 01:55:16 2015, ericq, Update, LSC, Uncooperative AUX X 
    Reply  Fri Dec 18 17:37:04 2015, rana, Update, LSC, Uncooperative AUX X 
       Reply  Mon Dec 21 02:29:49 2015, ericq, Update, LSC, AUX X RIN measurements auxXRIN.pdf
          Reply  Tue Jan 5 02:54:38 2016, ericq, Update, LSC, AUX X Freq Noise attempt 
             Reply  Tue Jan 5 13:17:06 2016, rana, Update, LSC, AUX X Freq Noise attempt 
                Reply  Tue Jan 5 16:33:45 2016, ericq, Update, LSC, AUX X Freq Noise attempt 
                   Reply  Thu Jan 7 04:28:39 2016, ericq, Update, LSC, AUX X Freq Noise measured PLLspec.pdfPLL_OLG.pdf
                      Reply  Thu Jan 7 16:52:32 2016, ericq, Update, LSC, AUX X Freq Noise measured AUXfreqnoise.pdfauxXmeasurements.zip
                         Reply  Thu Jan 7 19:04:25 2016, Koji, Update, LSC, AUX X Freq Noise measured 
                            Reply  Fri Jan 8 14:47:33 2016, ericq, Update, LSC, AUX Y Freq Noise measured AUX_freqnoises.pdfAUXY_Jan8.zip
                               Reply  Fri Jan 8 20:02:49 2016, rana, Update, LSC, AUX Y Freq Noise measured 
                               Reply  Sat Jan 9 00:39:15 2016, gautam, Update, LSC, AUX Y Freq Noise re-measured AuxPLL.pdf
                                  Reply  Mon Jan 11 19:01:56 2016, gautam, Update, LSC, PLL Marconi Investigation 6x
                                     Reply  Tue Jan 12 03:03:55 2016, ericq, Update, LSC, Frequently making noise 
                                        Reply  Tue Jan 12 19:38:31 2016, gautam, Update, LSC, Frequently making noise 2016_01_AUXLaser.tar.gzOLGs.pdfvariedCurrent.pdfvariedTemp.pdfPLL_setup.pdf
Message ID: 11908     Entry time: Tue Jan 5 02:54:38 2016     In reply to: 11894     Reply to this: 11910
Author: ericq 
Type: Update 
Category: LSC 
Subject: AUX X Freq Noise attempt 

[ericq, Gautam]

We set out to lock a marconi to the IR fiber beat of PSL + AUX X to measure some frequency noise, and failed.

In short, the Marconi's 1.6MHz max external FM isn't enough oomph to stabilize the PLL error signal. It's actually evident on the Agilent that the beat moves around a few times more than that, which I should've noticed sooner... We could briefly "lock" the PLL for a few tenths of a second, but weren't able to get a spectrum from this.

We also tried using the digital phase tracker temperature servo for some help at ~DC; this worked to the extent that we didn't have to twiddle the Marconi carrier frequency to stay on top of the fringes as the beat wandered, but it didn't otherwise stabilize the beat enough to make a difference in locking the PLL.

I suppose one more thing to try is to lock the PSL laser itself to each AUX laser in turn via PLL, and look for different / excess noise.

The Green and IR beat electronics are a in a little bit of disarray at the moment, but it's not like anyone else is going to be using them for the time being...

ELOG V3.1.3-